OpenStudy (hellomynameiskam):

What is the product in lowest terms? - 8/9 x 5/6 A. –1 B.- 13/15 C.- 20/27 D. - 1/5

OpenStudy (jameshorton):

1/4 * 3/5 Multiply the numerators and denominators together: 1 * 3 = 3 4 * 5 = 20 3/20 3/4 * 5/9 Multiply the numerators and denominators together: 3 * 5 = 15 4 * 9 = 36 15/36 Divide both by 3: 15 / 3 = 5 36 / 3 = 12 5/12 Comments Report 1

OpenStudy (jameshorton):

5/12

OpenStudy (hellomynameiskam):

Theres no 5/12 as an option though.

OpenStudy (hellomynameiskam):

@jameshorton

OpenStudy (hellomynameiskam):

Thats not even the right problem...